OpenStudy (asanchez):

Parallelogram LONM ON=2M+8 OL=S+1 LM= 3M-1 NM=5M-2 It says to find the values of the variables, the find the lengths of the sides.

OpenStudy (anonymous):

LM=ON, so 3m-1=2m+8, so m=9. in which case it is 26 by 43. but are you sure you typed it correctly? look at OL again

OpenStudy (asanchez):

yeah it says OL= S+1 SO IDK.

OpenStudy (anonymous):

and NM=5m-2? they have diff variables?

OpenStudy (asanchez):

Yes, thats why i got confused.

OpenStudy (anonymous):

well then i guess S=42, to make S+1=43

OpenStudy (asanchez):

Thanks again. (: